OpenStudy (anonymous):

Convert this percent to a decimal. 9 10 × 1% = _____ Select one of the options below as your answer: A. .00009 B. .0009 C. .009 D. .09

OpenStudy (jiteshmeghwal9):



OpenStudy (anonymous):

Show Me How To Do It

OpenStudy (anonymous):

@jiteshmeghwal9 I Cant Read That

OpenStudy (anonymous):

9/10 = 0.9 1% = 1/100 = 0.01 So 9/10 * 1% would be 0.9 * 0.01 which is 0.09.

OpenStudy (jiteshmeghwal9):

\[\frac{9}{10} \times \frac{1}{100}=0.9 \times 0.01=?\]@DamnImBad :)

OpenStudy (anonymous):

c